6. Multiply the complex numbers, write the product in the form a + bi a. (5 + 3i)(5-3i) b. (5 + 2i)(7-1)
Answer #1

Answer

1> 34 ; 2> 37 + 9i

1> ( 5 + 3i ) ( 5 - 3i )

we know that ,   

( a+b ) ( a-b ) = a2 - b2

i 2 = -1

so

( 5 + 3i ) ( 5 - 3i ) = (52) - (3 i)2 = 25 - (32)(i2) = 25 - 9(-1) = 25 + 9 = 34

2 > ( 5 + 2i ) (7 - i ) = 5( 7 - i ) + 2i( 7 - i )

= ( 5x7 - 5i ) + ( (2i)7 - (2i)i ) = ( 35 - 5i ) + (14i - (2i2))

= 35 - 5i + 14i - 2 (-1) = 35 + 9i + 2 = 37 + 9i

pls Upvote ANSWER

THANKYOU